# LaCo: Large Language Model Pruning via Layer Collapse

### v1

@@ 4.1.1 Post-training
Due to the inevitable performance loss caused by pruning, we investigate whether models pruned using our LaCo can effectively inherit parameters from the original model and quickly recover performance through post-training. Specifically, we select the pruned Llama2-7B and Baichuan2-7B models obtained through LaCo in the main experiments and post-train them. For training pruned Llama2-7B, we utilize approximately 1.0 billion tokens from the English dataset, while for pruned Baichuan2-7B, we employ approximately 1.25 billion tokens, with a 50% from English and Chinese. The detailed implementation can be found in the Appendix B.
@@
We also evaluate the post-trained models on multiple benchmarks, and the detailed results can be found in Appendix C Table 9. We list the average score for each category and the overall average score in Table 3.
@@
Table 3: The average scores across all categories and the overall average score of pruned models, post-trained models, as well as post-trained models followed by re-pruning.
@@ 4.1.2 Re-pruning
Since it is possible to partially restore performance using post-training on an LLM with approximately 25%-30% of its parameters pruned, it raises the question of whether we can further prune the post-trained model to obtain one with only around 50% of the original parameters while still maintaining relatively good model performance. To this end, we further prune the previously post-trained pruned Llama2-7B model using LaCo, resulting in a model with 17 layers, retaining 55% of the parameters of the original Llama2-7B model with 32 layers. We evaluate the re-pruned model. The detailed results are shown in Appendix C Table 9 and the average results are in Table 3.
From the tables, it can be seen that although the model has only 55% of the parameters of the original model, it still retains an average of approximately 70% of the original 7B model performance. However, our training data quality and scale are limited. With access to more and better training data, our LaCo method should be able to demonstrate even greater utility in this pruning+post-training+re-pruning pipeline on larger models.
@@ 4.2 Layer-wise Similarity
This section discusses our motivation, i.e., why we consider merging adjacent layers. Our motivation primarily stems from the observation that the changes in parameters and output representations between adjacent layers within the LLM are not particularly significant.
In Figure 4, we illustrate the L2 similarities between the SAN q,k,v matrices of each layer and their corresponding matrices from the subsequent one layer as well as the up and down matrices of the MLP for both Llama2-7B and Baichuan2-7B. The results indicate that the maximum L2 values between corresponding matrices in adjacent layers are generally no more than 200. However, the sizes of the up and down matrices of the MLP reach 11008*4096, and the SAN q,k,v matrices also contain 4096*4096 parameters. Therefore, the change in each element of the matrices between adjacent layers is very small.
In Figure 5 (a), we randomly select 20 sentences from Wikipedia and calculate the cosine similarity between the hidden-states of adjacent layers outputs. The results show that for both Baichuan2-7B and Llama2-7B, the representation similarity between adjacent layers from layers 3 to 28 is typically very close to 1, which suggests that the representations between adjacent layers of LLMs also exhibit high similarity. Therefore, the high similarity in parameters and representations between adjacent layers leads us to consider whether a single layer can replace multiple layers following it.
Moreover, the similarity in parameters suggests that we should pay more attention to the differences between layers. Inspired by previous model merging work Yu et al. (2023); Matena and Raffel (2022), we come up with collecting the parameter differences between layers and then merging them into a single layer.
To verify that RDSC Layer Merge can indeed replace multiple layers with a merged single layer, we conduct the following experiments. Within layers 10 to 19, we attempt to merge consecutive 4 layers into one layer sequentially, then evaluate the cosine similarity between the output of the merged layer and the output of the original last layer among those 4 layers, as shown in Figure 5 (b). The results showcase that the lowest cosine similarity on the 4096-dimensional vectors is above 0.996, confirming the effectiveness of the proposed RDSC Layer Merge in preserving representations.
@@ 4.3 Varying Pruning Ratio
Table 4: The model performance at different pruning ratios.
In this section, we explore the performance of LaCo at different pruning ratios. We conduct experiments on Llama2-7B and Llama2-13B, controlling the pruning ratios at approximately 10%, 25% (our main experiments), and around 50% by setting different hyperparameters (as shown in Appendix A Table 6). We evaluate pruned models accordingly. The average results are shown in Table 4 and the detailed results are shown in Appendix C Table 10.
From the results, it can be observed that as the pruning ratio increases, there is an overall decrease in model performance. However, from around 10-15% pruning ratio to around 25% pruning ratio, the model performance does not decrease significantly, which indicates that our method remains relatively stable within this range. Furthermore, at close to 50% pruning ratio, the model still maintains approximately 70% of its performance, which demonstrates that our method does not cause the model to crash even when about half of its parameters are removed.
@@ 5 Related Work
How to reduce the excessively high training and inference costs caused by the increase in the size of LLMs has been a critical issue of concern. In recent years, the methods to alleviate this issue have mainly consisted of model quantization, knowledge distillation, and model pruning.
Model Quantization reduces the size of models by converting their weights from high-precision floating-point representations to lower-precision floating-point or integer representations. SmoothQuant Xiao et al. (2023) quantizes both weights and activation values and smooths the activation outliers. Gptq Frantar et al. (2022) leverages approximate second-order information to weight quantization. Qlora Dettmers et al. (2023a) backpropagates gradients through a frozen, 4-bit quantized model into Low Rank Adapters. OmniQuant Shao et al. (2023) quantizes LLMs by optimizing various quantization parameters.
@@
Model Pruning refers to the technique for improving the efficiency of models by sparsification or removing parameters. Non-structured pruning often involves model sparsity. SparseGPT Frantar and Alistarh (2023) works by reducing the pruning problem to a set of extremely large-scale instances of sparse regression. SpQR Dettmers et al. (2023b) involves the identification and isolation of outlier weights during LLM sparsification. Structured pruning primarily discards portions of model modules. LLM-Pruner Ma et al. (2023) selectively removes non-critical structures based on gradient information. ShearedLLaMA Xia et al. (2023) employs targeted structured pruning and dynamic batch loading to prune the Llama2.
However, model quantization and sparsification methods have a significant impact on model performance and typically require special hardware support. Knowledge distillation entails considerable training costs and is usually task-specific. Existing structured model pruning methods often disrupt the inherent structure of the model. In contrast, our LaCo does not disrupt the model structure during pruning and is also more concise, while maintaining excellent performance.
@@ 6 Conclusion
In this paper, we propose a concise layer-wise structured pruning method called Layer Collapse, which involves the collapse of rear model layers into the preceding layer, facilitating a swift reduction in model size. Our LaCo does not require special hardware support and preserves the intrinsic structure of the model. Experimental results demonstrate that the proposed LaCo significantly outperforms existing SOTA structured pruning methods. Subsequently, we conduct post-training on the pruned models, verifying that LaCo effectively inherits the parameters of the original model and enables rapid convergence of the pruned model during post-training. Later, we also discuss our motivation from the perspective of layer-wise similarity. Finally, we explore the performance of models pruned by LaCo at different pruning ratios.
@@ Appendix A Hyperparameter Settings
Table 6: Hyperparameter settings for varying pruning ratios.
Appendix B Post-Training Implementation Details
@@
For post-training, we utilize the code framework provided by the LLaMA-Factory repository hiyouga (2023) along with DeepSpeed ZeRO-2. The sequence length is set to 4096 following the default settings of the Llama2-7B and Baichuan2-7B. We use Adam optimizer with the learning rate of 2e-4, where we set the Adam parameters $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The batch size is set to 8 per GPU, resulting in a total batch size of 32. We configure the gradient accumulation steps to be 4 and utilize a cosine learning rate scheduler. Additionally, we apply weight decay of 0.1 and set the maximum gradient normalization to 1.0.
Appendix C Supplementary Results

### v2

@@ 4.1.1 Post-training
Due to the inevitable performance loss caused by pruning, we investigate whether models pruned using our LaCo can effectively inherit parameters from the original model and quickly recover performance through post-training on the full parameters. Specifically, we select the pruned Llama2-7B and Baichuan2-7B models obtained through LaCo in the main experiments and post-train them. For training pruned Llama2-7B, we utilize approximately 1.0 billion tokens from the English dataset, while for pruned Baichuan2-7B, we employ approximately 1.25 billion tokens, with a 50% from English and Chinese. The detailed implementation can be found in the Appendix C.
@@
We also evaluate the post-trained models on multiple benchmarks with detailed results in Appendix E Table 24. The average scores for each category and the overall average are in Table 6.
@@
Table 6: Average scores across all categories and the overall average score of pruned models, post-trained models, post-trained models followed by re-pruning.
@@
LaCo achieves excellent performance through post-training, prompting us to compare its effectiveness with the SOTA LLM-Pruner on the same training data. Our results, shown in the Appendix D.3 Table 21, indicate that the model pruned by LaCo outperforms the one pruned by LLM-Pruner after post-training. Meanwhile, LaCo also significantly reduces training resource consumption.
@@ 4.1.2 Re-pruning
Since it is possible to partially restore performance using post-training on an LLM with approximately 25%-30% of its parameters pruned, it raises the question of whether we can further prune the post-trained model to obtain one with only around 50% parameters while still maintaining relatively good performance. Thus, we further prune the previously post-trained pruned Llama2-7B model using LaCo, resulting in a model with 17 layers, retaining 55% of the parameters of the original Llama2-7B model. We evaluate the re-pruned model. The detailed results are shown in Appendix E Table 24 and the average results are in Table 6.
The tables show that even with only 55% parameters, the model still retains about 70% of the original 7B model performance. However, our training data quality and scale are limited. With more and better training data, LaCo should demonstrate even greater utility in the pruning+post-training+re-pruning pipeline on larger models.
@@ 4.2 Layer-wise Similarity
This section discusses our motivation for merging adjacent layers. Our primary motivation comes from observing that the changes in parameters and output representations between adjacent layers in the LLMs are not particularly significant.
In Figure 4, we show the L2 similarities between the SAN q, k, v matrices of each layer and their counterparts in the subsequent layer, as well as the upscaling and downscaling matrices of the MLP for both Llama2-7B and Baichuan2-7B. The results indicate that the maximum L2 values between corresponding matrices in adjacent layers are generally no more than 200. Given the large sizes of the MLP upscaling (11008x4096) and SAN q, k, v (4096x4096) matrices, the change in each element between adjacent layers is minimal.
In Figure 5 (a), we randomly select 20 sentences from Wikipedia and calculate the cosine similarity between the hidden-states of adjacent layers outputs. The results show that for both Baichuan2-7B and Llama2-7B, the representation similarity between adjacent layers from layers 3 to 28 is typically very close to 1. The high similarity in parameters and representations between adjacent layers leads us to consider that a single layer might replace multiple subsequent layers.
Moreover, the similarity in parameters suggests focusing on the differences between layers. Inspired by previous model merging work Yu et al. (2023); Matena and Raffel (2022), we come up with collecting parameter differences between layers and merging them into a single layer. To verify that RDSC Layer Merge can replace multiple layers with one, we conduct the experiment: we merge every four consecutive layers into one within layers 10 to 19 and evaluate the cosine similarity between the merged layer’s output and the original last layer’s output, as in Figure 5 (b), where the lowest cosine similarity on the 4096-dimensional vectors is above 0.996, confirming the effectiveness of RDSC Layer Merge in preserving representations.
@@ 4.3 Varying Pruning Ratio
Table 7: Model performance at different pruning ratios.
In this section, we explore the performance of LaCo at different pruning ratios. We conduct experiments on Llama2-7B and Llama2-13B, controlling the pruning ratios at approximately 10%, 25% (our main experiments), and around 50% by setting different hyperparameters (as shown in Appendix A Table 9) [^3] Further ablation study on the hyperparameters $\mathcal{C}$, $\mathcal{I}$, $\mathcal{D}$, $\mathcal{T}$, different similarity metrics, different merging strategies, different calibration datasets can be found in Appendix B.. We evaluate pruned models accordingly. The average results are shown in Table 7 and the detailed results are shown in Appendix E Table 25.
As the pruning ratio increases, overall model performance decreases. However, from a pruning ratio of around 10-15% to about 25%, the performance does not significantly decline, indicating our method’s stability within this range. Furthermore, at a pruning ratio close to 50%, the model still maintains approximately 70% performance, demonstrating that our method prevents model crashes even with about half of the parameters removed.
@@ 5 Related Work
Model Quantization reduces model size by converting weights from high-precision floating points to lower-precision floating points or integers. SmoothQuant Xiao et al. (2023) quantizes both weights and activations while smoothing activation outliers. GPTQ Frantar et al. (2022) uses approximate second-order information for quantization. QLoRA Dettmers et al. (2023a) backpropagates gradients through a frozen, 4-bit quantized model into Low Rank Adapters. OmniQuant Shao et al. (2023) optimizes various quantization parameters.
@@
Model Pruning refers to techniques for improving model efficiency by sparsification or parameter removal. Non-structured pruning often involves model sparsity. SparseGPT Frantar and Alistarh (2023) reduces the pruning problem to large-scale instances of sparse regression, while SpQR Dettmers et al. (2023b) identifies and isolates outlier weights during LLM sparsification. Structured pruning primarily removes parts of model modules. LLM-Pruner Ma et al. (2023) selectively eliminates non-critical structures based on gradient information. DirectShare Cao et al. (2024) uses activation sharing methods to improve LLM inference efficiency. ShearedLLaMA Xia et al. (2023) uses targeted structured pruning and dynamic batch loading to prune Llama2. FLAP An et al. (2024) assesses the recoverability of the output feature map after weight removal using the fluctuation pruning metric and standardizes importance scores to adaptively define the global compressed model structure. Shortened LLaMA Kim et al. (2024) demonstrates that depth pruning can efficiently compress LLMs while matching or surpassing the performance of recent width pruning methods.
However, model quantization and sparsification typically require special hardware and usually impact performance. Knowledge distillation is costly and task-specific. Existing structured pruning methods often disrupt the model inherent structure. In contrast, LaCo maintains the model structure, which is more concise and preserves excellent performance. Although some existing works Din et al. (2023); Fan et al. (2019); Belrose et al. (2023) have utilized layer-skipping/dropping to accelerate inference, LaCo is fundamentally different. It is the first pruner based on layer collapse, resulting in a smaller, faster, more memory-efficient model with strong performance. Furthermore, those methods typically require training new parameters to determine which layers to skip/drop during inference, whereas LaCo does not require any training.
@@ 6 Conclusion
In this paper, we propose a concise layer-wise structured pruning method called Layer Collapse (LaCo), which merges rear model layers into preceding layers for rapid model size reduction. LaCo does not require special hardware support and preserves the model’s intrinsic structure. Experimental results show that LaCo significantly outperforms current SOTA structured pruning methods, also revealing potential parameter redundancy in existing LLMs. We conduct ablation studies on various settings of LaCo. We also post-train the pruned models, confirming that LaCo effectively inherits the original model parameters. Additionally, we discuss our motivation from the perspective of layer-wise similarity and explore the performance of LaCo-pruned models at different pruning ratios.
@@
Our motivation comes from current model merging techniques, but like existing baselines (LLM-Pruner Ma et al. (2023) and SliceGPT Ashkboos et al. (2024)), our method lacks a complete theoretical proof. We consider this as future work.
Additionally, there may be better merging methods, even though our experimental results demonstrate that LaCo’s current merging approach is effective. We will continue to search for improved layer merging methods in the future.
@@ Appendix A Hyperparameter Settings
Table 9: Hyperparameters for varying pruning ratios.
Appendix B Ablation Study
B.1 Impact of Number of Layers Merged per Operation $\mathcal{C}$
We conduct an ablation study on the number of layers to be merged during each merging operation $\mathcal{C}$, as it is one of the key parameters that control the compression rate. The results are shown in Table 10, where the experiment is based on the Llama2-7b with a 27% compression rate.
BoolQ PIQA HeSw
$\mathcal{C}=1$ (Result in Table 15) 52.57 68.17 48.61
$\mathcal{C}=2$ 63.12 68.93 54.98
$\mathcal{C}=3$ 63.98 69.71 55.34
$\mathcal{C}=4$ (Result in Table 1) 64.07 69.80 55.69
$\mathcal{C}=5$ 64.15 70.02 55.76
$\mathcal{C}=6$ 64.00 69.89 55.49
Table 10: Ablation study on $\mathcal{C}$.
When $\mathcal{C}$ is too small, the model’s performance degrades. When $\mathcal{C}$ is between 3 and 6, the performance remains relatively good. If $\mathcal{C}$ is too large, it causes the cosine similarity of the hidden-states to drop quickly below the threshold, stopping the pruning loop prematurely and resulting in a lower compression rate. Therefore, we recommend setting $\mathcal{C}$ between 4 and 7.
B.2 Impact of Minimum Interval $\mathcal{I}$
To explore the impact of different values of $\mathcal{I}$, we also conduct an ablation study. The experiment is based on the Llama2-7b model with a 27% sparsity. The results are shown in Table 11.
BoolQ PIQA HeSw
$\mathcal{I}=1$ 63.99 68.90 55.79
$\mathcal{I}=2$ (Result in Table 1) 64.07 69.80 55.69
$\mathcal{I}=3$ 64.05 69.57 55.48
$\mathcal{I}=4$ 62.29 68.92 54.87
Table 11: Ablation study on $\mathcal{I}$.
The results indicate that the model performs well when $\mathcal{I}$ is between 1 and 3. However, when $\mathcal{I}$ is set to 4, the model’s performance declines. This observation aligns with our empirical findings: modifying the parameters of layers closer to the input end (such as layers 1-12) can lead to performance drops. Setting $\mathcal{I}$ to 4 causes the merge operations to occur closer to the input end. Therefore, we choose $\mathcal{I}=2$, as it relatively maintains better model performance.
B.3 Impact of Dataset $\mathcal{D}$
To understand the impact of different datasets $\mathcal{D}$ on LaCo’s effectiveness, we conduct an ablation study using Llama2-7B. We perform three rounds of pruning, each time selecting different sets of 10 sentences as $\mathcal{D}$, with a 27% compression ratio to match Table 1. The results are shown in Table 12.
BoolQ PIQA HeSw
Result in Table 1 64.07 69.80 55.69
Round1 64.50 70.28 56.14
Round2 63.70 69.14 54.92
Round3 63.90 69.66 55.18
Table 12: Ablation study on $\mathcal{D}$, where the model is Llama2-7B with compression ratio of 27%.
We can find that the results are all on nearly the same scale, indicating that the random selection of sentences in $\mathcal{D}$ has no notable impact on the results.
B.4 Impact of Threshold $\mathcal{T}$
We also aim to understand the impact of threshold $\mathcal{T}$ on the model pruned by LaCo. For the results of Llama2-7B in Table 1, we keep other parameters unchanged and set $\mathcal{T}$ to 0.85, 0.45, and 0.25 ($\mathcal{T}=0.65$ corresponds to the results in Table 1). We evaluate the pruned models on several datasets. The results, shown in Table 13, indicate that a smaller threshold leads to a larger compression ratio and poorer performance, which aligns with intuition.
Ratio/Layer BoolQ PIQA HeSw
$\mathcal{T}=0.65$ 27.1%/23 64.07 69.80 55.69
$\mathcal{T}=0.85$ 9.0%/29 70.92 76.01 68.15
$\mathcal{T}=0.45$ 48.0%/16 59.82 60.34 36.09
$\mathcal{T}=0.25$ 60.1%/12 41.68 51.74 26.37
Table 13: Ablation study on $\mathcal{T}$. The model is Llama2-7B. $\mathcal{T}=0.65$ corresponds to the results in Table 1.
B.5 Different Similarity Metrics
we use the cosine similarity of the representations from the final layer outputs as a metric for LaCo. We also aim to explore the feasibility of using common distribution distances as metrics, such as KL divergence and kernel/linear CKA (Centered Kernel Alignment Kornblith et al. (2019)). Specifically, we replace cosine similarity with distribution distances to measure the difference in representations. However, we find that the distribution distances almost remain constant, as in Table 14.
KL Divergence Kernel CKA Linear CKA
Constant 0.00 1.00 1.00
Table 14: The distribution distances of model output representations.
KL Divergence being nearly 0 and CKA being nearly 1 indicate minimal differences in the distributions of the final layer outputs. This suggests that the LaCo merging process does not significantly alter the model output. Additionally, due to the high dimensionality of the representations, the distribution distances tend to be constant, resulting in a lack of discrimination. Therefore, we chose the more discriminative cosine similarity.
B.6 Drop or Merge
As shown in Eq. 1, LaCo merges multiple adjacent layers into one. This leads us to consider an extreme case: if we set $m=1$, LaCo will no longer merge layers but simply drop a layer. We also aim to explore the performance differences between the drop and merge operations in LaCo. Thus, we conduct an experiment with $m=1$ on Llama2-7B, setting the compression rate to the same 27% as in Table 1. The results on some benchmarks are shown in Table 15.
BoolQ PIQA HeSw
Drop 52.57 68.17 48.61
Mege (Result in Table 1) 64.07 69.80 55.69
Table 15: Results of the drop or merge operation.
The results indicate that the drop operation is not as effective as the merge operation and $m=1$ is not a good hyperparameter setting. The merge operation aligns better with our intention.
B.7 Iterative-based or Rule-based Merge
LLM Strategy BoolQ PIQA HeSw
Llama2-7B Rule 63.49 68.72 53.16
LaCo 64.07 69.80 55.69
Baichuan2-7B Rule 52.94 67.28 48.78
LaCo 56.15 68.50 52.28
Table 16: The results of using rule-based merging and LaCo iterative merging.
We want to determine if our iterative search-based merging strategy is superior to rule-based merging. To test this, we perform rule-based merging on Llama2-7B and Baichuan2-7B, both with 32 layers. We merge layers in groups of four, starting from the top, specifically merging layers (29, 30, 31, 32), (21, 22, 23, 24), and (13, 14, 15, 16). We avoid merging before the 16th layer due to significant performance drops observed in those cases. The resulting models achieved compression rates equivalent to those in Table 1. The results in Table 16 indicate that LaCo performs better than the rule-based approach. Notably, even simple rule-based merging can outperform baselines across multiple datasets, demonstrating the potential of merging for model compression.
B.8 Effect of Calibration Dataset from Different Sources
We explore whether calibration datasets from different sources will affect the performance.
Specifically, we select 10 sentences from the BookCorpus Kobayashi (2018) dataset used by LLM-Pruner as the calibration dataset. We prune Llama2-7B at a 27% compression rate and evaluate the results on several datasets. We repeat the experiments three times.
BoolQ PIQA HeSw
Wikipedia 64.07 69.80 55.69
BookCorpus-Round1 64.15 69.60 55.87
BookCorpus-Round2 63.80 69.45 55.61
BookCorpus-Round3 64.30 70.08 55.98
Table 17: Results of calibration datasets from different sources.
The results in Table 17 are at the same level as those in our paper. This demonstrates that, given high-quality corpora, different calibration datasets have little impact on the performance of the models pruned with LaCo at the same compression rate.
B.9 Effect of Calibration Dataset Size
We want to explore how the size of the calibration dataset affects the performance of LaCo. Specifically, we take 5, 10, 50, 100, and 200 samples from Wikipedia and prune Llama2-7B at a 27% compression rate. The results are shown in the Table 18, and it can be seen that the size of the calibration dataset has little impact on the results.
BoolQ PIQA HeSw
Result in Table 1 (10 samples) 64.07 69.80 55.69
5 samples 63.89 69.74 55.54
50 samples 64.10 69.69 55.81
100 samples 64.25 69.81 55.59
200 samples 63.98 69.92 55.62
Table 18: Results of different calibration dataset size.
Therefore, we chose a small calibration dataset to ensure both the pruning effect and the pruning speed. The size of our calibration dataset is the same as that used by the LLM-Pruner, which is 10 samples.
Appendix C Post-Training Implementation Details
@@
We use the LLaMA-Factory Zheng et al. (2024) framework along with DeepSpeed ZeRO-2. The sequence length is set to 4096, following the default settings for Llama2-7B and Baichuan2-7B. We use the Adam optimizer with a learning rate of 2e-4, setting $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The batch size is 8 per GPU, resulting in a total batch size of 32, with gradient accumulation steps set to 4. We employ a cosine learning rate scheduler, apply a weight decay of 0.1, and set the maximum gradient normalization to 1.0.
Appendix D Supplementary Results (Part 1)
D.1 Results on Llama2-70B
We supplement the results using Llama2-70B with a compression rate set to 30%. We conduct experiments on CMNLI, HeSw, PIQA, and BoolQ, and Table 19 presents the average results.
Pruner Dense LaCo LLM-Pruner SliceGPT
Average 65.92 57.91 53.78 51.36
Table 19: The average scores on several benchmarks using different pruners with a 30% compression rate on Llama2-70B.
The results demonstrate that our method can also scale up effectively. It performs better than the baselines even on the 70B model.
D.2 Memory Consumption and Inference Speed
LLM Pruner Memory (MB) Infer. (tokens/s)
Llama2 -7B Dense 13410 38.53
LLMPru. 10434 33.22 ($\downarrow$)
SliceGPT 11770 44.88
LaCo 9894 50.80
Llama2 -13B Dense 25902 29.98
LLMPru. 19874 27.15 ($\downarrow$)
SliceGPT 22506 35.16
LaCo 19422 38.65
Baic2. -7B Dense 14810 37.13
LLMPru. 11898 38.95 ($\downarrow$)
SliceGPT 13586 36.67
LaCo 11716 49.15
Baic2. -13B Dense 27410 36.93
LLMPru. 22390 31.61 ($\downarrow$)
SliceGPT 23956 29.35 ($\downarrow$)
LaCo 21010 47.46
Table 20: The memory consumption and average inference speed on English Wikipedia dataset for different pruned models. $\downarrow$ means the performance worse than the Dense model.
D.3 Comparison of Post-trained Pruned Models
GPU*hour BoolQ PIQA HeSw
LaCo 88 60.26 65.01 45.49
LLM-Pru. 216 58.75 61.26 43.53
Table 21: Results on different datasets for models pruned to 55% sparsity using LaCo and LLM-Pruner on Llama2-7B, followed by post-training on the same data.
We prune Llama2-7B to 55% sparsity using LaCo and LLM-Pruner, then conduct post-training with the same data in Table 6. The results in Table 21 show that LaCo performs better. Additionally, training the model pruned by LLM-Pruner requires 216 GPU*hours (27 hours on 8 A100 GPUs), while only 88 GPU*hours (22 hours on 4 A100 GPUs) for LaCo-pruned model. Thus, LaCo saves more computational resources for post-training.
D.4 Performance Percentage for Main Results